# Anatomy of an ionized bubble: NIRCam grism spectroscopy of the double-peaked Lyman- emitter COLA1 and its environment

### v1

@@ Acknowledgements.
- Calabro et al. (2024) Calabro, A., Castellano, M., Zavala, J. A., et al. 2024, arXiv e-prints, arXiv:2403.12683
@@
- Chen et al. (2023) Chen, Z., Stark, D. P., Mason, C., et al. 2023, arXiv e-prints, arXiv:2311.13683
@@
- Choustikov et al. (2023) Choustikov, N., Katz, H., Saxena, A., et al. 2023, arXiv e-prints, arXiv:2304.08526
- Choustikov et al. (2024) Choustikov, N., Katz, H., Saxena, A., et al. 2024, arXiv e-prints, arXiv:2401.09557
@@
- Dayal & Giri (2023) Dayal, P. & Giri, S. K. 2023, arXiv e-prints, arXiv:2303.14239
@@
- Endsley et al. (2023) Endsley, R., Stark, D. P., Whitler, L., et al. 2023, arXiv e-prints, arXiv:2306.05295
@@
- Ferrara (2023) Ferrara, A. 2023, arXiv e-prints, arXiv:2310.12197
@@
- Greene et al. (2023) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2023, arXiv e-prints, arXiv:2309.05714
@@
- Jung et al. (2023) Jung, I., Finkelstein, S. L., Arrabal Haro, P., et al. 2023, arXiv e-prints, arXiv:2304.05385
@@
- Kerutt et al. (2023) Kerutt, J., Oesch, P. A., Wisotzki, L., et al. 2023, arXiv e-prints, arXiv:2312.08791
@@
- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, arXiv e-prints, arXiv:2401.09981
@@
- Leclercq et al. (2024) Leclercq, F., Chisholm, J., King, W., et al. 2024, arXiv e-prints, arXiv:2401.14981
@@
- Lu et al. (2023) Lu, T.-Y., Mason, C., Hutter, A., et al. 2023, arXiv e-prints, arXiv:2304.11192
@@
- Mascia et al. (2023a) Mascia, S., Pentericci, L., Calabrò, A., et al. 2023a, arXiv e-prints, arXiv:2309.02219
- Mascia et al. (2023b) Mascia, S., Pentericci, L., Calabrò, A., et al. 2023b, A&A, 672, A155
- Mascia et al. (2023c) Mascia, S., Pentericci, L., Saxena, A., et al. 2023c, A&A, 674, A221
@@
- Matthee et al. (2023a) Matthee, J., Mackenzie, R., Simcoe, R. A., et al. 2023a, ApJ, 950, 67
- Matthee et al. (2023b) Matthee, J., Naidu, R. P., Brammer, G., et al. 2023b, arXiv e-prints, arXiv:2306.05448
@@
- Naidu et al. (2022b) Naidu, R. P., Oesch, P. A., Setton, D. J., et al. 2022b, arXiv e-prints, arXiv:2208.02794
@@
- Nakane et al. (2023) Nakane, M., Ouchi, M., Nakajima, K., et al. 2023, arXiv e-prints, arXiv:2312.06804
@@
- Saxena et al. (2023a) Saxena, A., Bunker, A. J., Jones, G. C., et al. 2023a, arXiv e-prints, arXiv:2306.04536
@@
- Saxena et al. (2023b) Saxena, A., Robertson, B. E., Bunker, A. J., et al. 2023b, A&A, 678, A68
@@
- Topping et al. (2023) Topping, M. W., Stark, D. P., Endsley, R., et al. 2023, arXiv e-prints, arXiv:2307.08835
- Topping et al. (2024) Topping, M. W., Stark, D. P., Senchyna, P., et al. 2024, arXiv e-prints, arXiv:2401.08764
@@
- Übler et al. (2023) Übler, H., Maiolino, R., Pérez-González, P. G., et al. 2023, arXiv e-prints, arXiv:2312.03589
@@
- Whitler et al. (2023) Whitler, L., Stark, D. P., Endsley, R., et al. 2023, arXiv e-prints, arXiv:2305.16670
@@
- Witstok et al. (2023) Witstok, J., Smit, R., Saxena, A., et al. 2023, arXiv e-prints, arXiv:2306.04627
@@ Appendix A C1F [O III] emitter sample catalog
In Table 4 we list the C1F sample, composed of 141 [O III] emitters, selected as described in Sect. 3.1. We tabulate the IDs, sky coordinates of the F356W detection in the J2000 system, [O III] doublet redshifts (see Sect. 3.2), Prospector UV magnitudes (see Sect. 3.3) and [O III]5008 luminosities measured from the grism data.
@@ Appendix C Examples of spectra
In Figures 15 and 16 we show examples of inspection images of selected [O III] emitters in this work. The first five objects shown in Fig. 15 correspond to those in the environment of COLA1 (). The object with ID NUMBER 1272 illustrates how the morphology of sources is captured in the emission lines of 2D grism spectra, in many cases allowing a secure match with the imaging. On the other hand, the object with ID NUMBER 18752 has a prominent extended morphology as well, and its [O III] doublet is only seen in module B. Since modules A and B disperse light in opposite directions, the spatial information is mirrored in the spectral direction for module B.

### v2

@@ Acknowledgements.
- Calabro et al. (2024) Calabro, A., Castellano, M., Zavala, J. A., et al. 2024, ApJ, submitted, arXiv:2403.12683
@@
- Chen et al. (2024) Chen, Z., Stark, D. P., Mason, C., et al. 2024, MNRAS, 528, 7052
@@
- Choustikov et al. (2024) Choustikov, N., Katz, H., Saxena, A., et al. 2024, MNRAS, 529, 3751
@@
- Dayal & Giri (2024) Dayal, P. & Giri, S. K. 2024, MNRAS, 528, 2784
@@
- Endsley et al. (2023a) Endsley, R., Stark, D. P., Whitler, L., et al. 2023a, MNRAS, 524, 2312
- Endsley et al. (2023b) Endsley, R., Stark, D. P., Whitler, L., et al. 2023b, MNRAS, submitted, arXiv:2306.05295
@@
- Ferland et al. (2013) Ferland, G. J., Porter, R. L., van Hoof, P. A. M., et al. 2013, Rev. Mexicana Astron. Astrofis., 49, 137
- Ferrara (2024) Ferrara, A. 2024, A&A, 684, A207
@@
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
@@
- Jung et al. (2024) Jung, I., Finkelstein, S. L., Arrabal Haro, P., et al. 2024, ApJ, 967, 73
@@
- Kerutt et al. (2024) Kerutt, J., Oesch, P. A., Wisotzki, L., et al. 2024, A&A, 684, A42
@@
- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, ApJ, 968, 38
@@
- Leclercq et al. (2024) Leclercq, F., Chisholm, J., King, W., et al. 2024, A&A, submitted, arXiv:2401.14981
@@
- Lu et al. (2023) Lu, T.-Y., Mason, C., Hutter, A., et al. 2023, MNRAS, submitted, arXiv:2304.11192
@@
- Mascia et al. (2024) Mascia, S., Pentericci, L., Calabrò, A., et al. 2024, A&A, 685, A3
- Mascia et al. (2023a) Mascia, S., Pentericci, L., Calabrò, A., et al. 2023a, A&A, 672, A155
- Mascia et al. (2023b) Mascia, S., Pentericci, L., Saxena, A., et al. 2023b, A&A, 674, A221
@@
- Matthee et al. (2023) Matthee, J., Mackenzie, R., Simcoe, R. A., et al. 2023, ApJ, 950, 67
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
@@
- Naidu et al. (2022b) Naidu, R. P., Oesch, P. A., Setton, D. J., et al. 2022b, ApJL, submitted, arXiv:2208.02794
@@
- Nakane et al. (2024) Nakane, M., Ouchi, M., Nakajima, K., et al. 2024, ApJ, 967, 28
@@
- Saxena et al. (2024) Saxena, A., Bunker, A. J., Jones, G. C., et al. 2024, A&A, 684, A84
@@
- Saxena et al. (2023) Saxena, A., Robertson, B. E., Bunker, A. J., et al. 2023, A&A, 678, A68
@@
- Topping et al. (2024a) Topping, M. W., Stark, D. P., Endsley, R., et al. 2024a, MNRAS, 529, 4087
- Topping et al. (2024b) Topping, M. W., Stark, D. P., Senchyna, P., et al. 2024b, MNRAS, 529, 3301
@@
- Übler et al. (2024) Übler, H., Maiolino, R., Pérez-González, P. G., et al. 2024, MNRAS, 531, 355
@@
- Whitler et al. (2024) Whitler, L., Stark, D. P., Endsley, R., et al. 2024, MNRAS, 529, 855
@@
- Witstok et al. (2024) Witstok, J., Smit, R., Saxena, A., et al. 2024, A&A, 682, A40
@@ Appendix A C1F [O III] emitter sample catalog
In Table 4 we list the C1F sample, composed of 141 [O III] emitters, selected as is described in Sect. 3.1. We tabulate the IDs, sky coordinates of the F356W detection in the J2000 system, [O III] doublet redshifts (see Sect. 3.2), Prospector UV magnitudes (see Sect. 3.4) and [O III]5008 luminosities measured from the grism data.
@@ Appendix C Examples of spectra
In Fig. 15 we show examples of inspection images of selected [O III] emitters in this work. The first five objects shown in Fig. 15 correspond to those in the environment of COLA1 (). The object with ID NUMBER 1272 illustrates how the morphology of sources is captured in the emission lines of 2D grism spectra, in many cases allowing a secure match with the imaging. On the other hand, the object with ID NUMBER 18752 has a prominent extended morphology as well, and its [O III] doublet is only seen in module B. Since modules A and B disperse light in opposite directions, the spatial information is mirrored in the spectral direction for module B.